### v1

# MapleGrasp: Mask-guided Feature Pooling for Language-driven Efficient Robotic Grasping

Abstract
[…]
Figure 1: Modular systems generate object masks and grasp poses, utilizing a control algorithm to estimate joint space trajectories, while VLAs directly map vision and textual modalities to actions.
1 Introduction
The ability to perceive and interact with objects in an environment is a defining characteristic of human intelligence. As humans mature, we develop capabilities to manipulate novel objects, generalizing our understanding beyond previous experiences. Replicating this intelligence in robotics remains a formidable challenge. Robots often encounter unfamiliar objects and must adapt to new scenarios without prior knowledge [29]. Foundational models, particularly Vision-Language Models (VLMs), aim to bridge this gap by leveraging extensive pre-trained multi-modal representations to enhance robotic perception and interaction [25, 64].
An end-to-end robotic manipulation framework typically comprises four key stages: task planning, visual grounding, object manipulation, and goal-oriented placement. Recent advances have demonstrated the effectiveness of VLMs across these stages, incorporating large-scale pre-trained representations to improve generalization [54]. An alternative paradigm involves Vision-Language-Action (VLA) models, which directly predict trajectories for grasping and manipulating objects, bypassing the modular pipeline [12, 4]. We highlight the difference in these two frameworks in Fig. 1. While VLAs have shown remarkable performance in long-range tasks across both simulation and real-world settings, their deployment in novel environments presents three key challenges: i) Task-Specific Data Collection: Collecting teleoperated or human demonstrations is often difficult and time-consuming for novel interactions [53]; ii) Expensive fine-tuning: Large-scale VLAs require task-specific fine-tuning, which is computationally expensive. Although parameter-efficient fine-tuning (PEFT) techniques alleviate some costs, they restrict adaptability by modifying only a small subset of parameters [20]; iii) Limited generalization: Fine-tuned VLAs often fail to generalize across dynamic environments, particularly when objects are repositioned or when visually similar distractor objects are introduced [46]. Given these challenges, modular frameworks offer an attractive alternative that balances generalizability with ease of deployment. By decoupling perception, planning, and manipulation, modular architectures allow state-of-the-art models in each component to be independently improved and integrated as research advances. Furthermore, LLM-agentic frameworks can serve as high-level decision-makers, dynamically invoking perception and manipulation modules for task execution in novel environments [15, 41, 42]. Compounded errors within modular frameworks can be mitigated using closed-loop feedback and error planning [6, 16].
In this work, we focus on a critical capability within modular frameworks, language-driven grasp detection. This task involves predicting a stable grasp pose for a specified object based on RGB-D images and natural language referring expressions. We focus on 2D grasp detection, which can be naturally extended to 3D using depth information and camera intrinsic parameters. Our contributions include:
  1. 1.
    A novel two-stage training strategy for LDRG: initially predicting text-referred object masks, then refining grasp predictions via mask-guided feature pooling.
  2. 2.
[…]
  3. 3.
    Introduce RefGraspNet, a large-scale open-source dataset comprising over 200 million grasp annotations, eight times larger than existing datasets, consisting of challenging real-world scenarios for grasping.
  4. 4.
    Extend 2D grasp predictions to 3D via depth maps and inverse kinematics, achieving performance comparable to recent Vision Language Action (VLA) models on the LIBERO benchmark, with superior generalization to diverse tasks while being parameter efficient.
2 Related Works
Grasp Detection. Robotic grasping with parallel grippers has been extensively studied in both the 2D and 3D domains. The 4-DoF grasp representation defines a grasp using four parameters: $(x,y)$ as the center of the grasp, $w$ as the grasp width, and $\theta$ as the grasp axis angle relative to a top-down hand orientation [1, 8, 17]. The 6-DoF representation extends this to 3D, incorporating three degrees of translation and three degrees of rotation required for stable object grasping [31, 10, 43]. Early approaches to object grasping relied on kinematics and geometric methods to identify stable grasp poses from depth maps [3] and 3D object models, using datasets generated through offline labeled training data, human demonstrations, or trial-and-error methods [13, 9]. The use of RGB-D data has shown promising results for precision grasping, where CNN and ResNet-based architectures leverage visual features to predict grasp pose parameters [22, 48, 65, 58]. Additionally, 3D-CNN models have been employed to extract 3D features from point clouds [39, 38, 51]. Recent advancements in transformer architectures and self-supervised learning have bridged the gap between vision and textual modalities, shifting research towards incorporating language as an additional modality for object manipulation [54]. Language-driven grasp detection (LDGD) refers to the task of associating referring text expressions with RGB-D images to predict a suitable robotic grasp pose for object manipulation [49]. Two primary approaches have emerged for LDGD: i) Visual grounding of referring expressions: This approach predicts segmentation masks for the target object [7, 55, 61], followed by the application of pre-trained grasp detection networks to compute grasp poses [2, 27, 32]. ii) End-to-end grasp pose estimation: This method directly predicts grasp poses from input text and RGB-D modalities [56, 63, 52].
Figure 2: Our framework trains on fused vision-language embeddings for object segmentation, then applies mask pooling in grasp prediction heads. MapleGrasp is an network-agnostic MFP block with no extra parameters and can be used with the CROG backbone as shown.
End-to-End Language-Driven Grasping. Our work focuses on utilizing RGB images of a scene along with free-form referring expressions to determine a stable 4-DoF grasp for the target object in cluttered environments. We extend this by incorporating depth maps to transform the 2D grasp into 3D and leveraging a differential inverse kinematics controller to compute the joint-space trajectory for grasp execution. Several studies have explored similar setups for LDGD [45, 46]. Recent work has introduced diffusion models that learn a probabilistic diffusion process to generate grasp poses during inference [49, 35, 36]. Vision-language fusion provides an alternative approach, offering faster convergence and lower inference times [34, 24, 54]. The ETRG framework enhances vision-language feature fusion for grasp detection by integrating vision-language adapters into frozen CLIP vision and text encoders [57]. MaskGrasp employs a tri-head architecture with cross-attention mechanisms using text embeddings, region-of-interest (ROI) feature embeddings, and mask embeddings [47]. GraspSAM adapts the Segment Anything Model [21] for grasp detection and utilizes GroundingDINO [28] to transform text expressions into object bounding boxes [37].
In our work, we analyze the impact of mask pooling on grasp detection and demonstrate that combining mask pooling with open-set object detectors significantly enhances performance on unseen objects—an issue that previous studies have not adequately addressed. Moreover, we present the first exhaustive comparison of LDRG approaches against large Vision-Language-Action (VLA) models for language-guided grasping tasks.
Vision-Language-Action Models. Direct modeling of vision and language modalities for joint-space robotic control has gained significant interest in recent research. Advances in vision-language models [59, 18, 30], coupled with the availability of large-scale robotic datasets spanning diverse environments, hardware platforms, and tasks [5], have enabled the training of Vision-Language-Action (VLA) models for robotic manipulation [40, 14, 26]. However, the large model sizes of VLAs often demand substantial GPU resources for fine-tuning and real-world deployment [50]. Recent research has sought to mitigate these computational challenges through parameter-efficient training techniques and action chunking methods [19, 20], enabling deployment on smaller GPUs with real-time kernel compatibility for seamless robotic control.
3 Mask-Guided Feature Pooling for Language Driven Grasp Detection
In this section, we introduce our approach of leveraging segmentation masks for feature pooling to focus on relevant regions in an image for targeted grasp pose estimation in cluttered environments. Furthermore, we present our two-stage training framework and discuss our improved training objective used to optimize our method.
3.1 Language-Driven Grasp Detection
Given an RGB-D image and a corresponding referring expression, language-driven grasping models predict a stable grasp pose to pick up the specified object using either a parallel gripper or a dexterous robotic hand. This task poses three primary challenges: i) accurately identifying the target object using textual cues provided by the referring expression, such as color, shape, and relative positioning; ii) determining a stable and collision-free grasp pose that can be reliably executed in real-world environments without interference from surrounding objects; and iii) generalizing grasp predictions effectively to previously unseen objects that were not included in the training dataset.
We focus on grasp pose estimation in 2D, where where grasps are typically represented as rectangles [49, 37, 34]. Each 2D grasp is defined by the center coordinates $(x,y)$ of the grasp rectangle, the grasp width $w$, and the orientation angle $\theta$ relative to a fixed horizontal axis. Given that this study focuses on top-down grasps performed with a parallel gripper, the height of the grasp rectangle remains constant. Such 2D grasp representations naturally extend into 3D space by incorporating depth information.
Formally, a 2D grasp $\mathbf{g_{2D}}=(x,\,y,\,w,\,\theta)$ is converted into a 3D grasp pose $\mathbf{T_{3D}}$ by projecting the pixel $(x,y)$ into 3D space. With depth value $D(x,y)$ at pixel $(x,y)$, the 3D coordinates relative to the camera are computed as:
$Z=D(x,y),\quad X=\frac{(x-c_{x})\,Z}{f_{x}},\quad Y=\frac{(y-c_{y})\,Z}{f_{y}},$ (1)
where $f_{x},f_{y}$ denote the camera focal lengths, and $c_{x},c_{y}$ represent the principal point offsets. The computed coordinates $(X,Y,Z)$ correspond to the grasp center position in the camera frame. The resulting top-down 3D grasp pose transformation matrix $\mathbf{T}_{3D}$ is then defined as:
$\mathbf{T}_{3D}=\begin{bmatrix}\cos\theta&-\sin\theta&0&X\\ \sin\theta&\cos\theta&0&Y\\ 0&0&1&Z\\ 0&0&0&1\end{bmatrix}.$ (2)
Direct prediction of all four grasp parameters (location, width, and orientation) at every pixel often yields suboptimal results due to their weak direct correlation with the original RGB-D data. Thus, contemporary approaches typically estimate grasp parameters using three distinct 2D prediction maps: i) Grasp Quality Map: Assigns a pixel-wise score for grasp quality; the pixel with the highest score represents the best grasp location, ii) Grasp Width Map: Encodes the optimal grasp width at each pixel and iii) Grasp Angle Map: Stores the grasp orientation $\theta$ at each pixel. This decoupled representation enhances the reliability and interpretability of grasp prediction models.
3.2 Referring Expression Segmentation
Language-Driven Grasp Detection (LDGD) shares strong similarities with a well-established computer vision task known as Referring Expression Segmentation (RES). The goal of RES is to predict a binary segmentation mask that separates the referred object from its surroundings [33, 62]. Since both tasks involve object-level visual grounding, several LDGD frameworks incorporate both segmentation mask prediction and grasp rectangle estimation from the referring query [47, 57]. These architectures typically contain separate projection layers for predicting the binary object mask and the grasp maps. However, since predictions are made independently, the segmentation mask may be correctly identified while grasp maps remain inaccurate.
3.3 Motivation for Mask Pooling
The predicted binary segmentation masks highlight the image regions corresponding to the referring expression, providing an implicit constraint for grasp pose estimation. If the segmentation mask is correctly predicted, then the optimal grasp pose should lie within the segmented region. This suggests that the grasp prediction module does not need to consider grasp locations outside the segmented object, provided that the segmentation branch can accurately localize the object. With recent advancements in VLMs for RES, zero-shot object segmentation has significantly improved. Consequently, off-the-shelf segmentation models can be used to extract masks for grasp pose estimation networks to predict stable grasp poses [7, 61]. We hypothesize that incorporating segmentation masks as a guiding mechanism within the grasp detection pipeline can improve accuracy by focusing predictions on relevant regions. Prior work has not explored this direction, and we argue that integrating mask pooling with LDRG can effectively utilize advances in RES to enhance grasp detection accuracy.
3.4 Mask-Guided Feature Pooling (MFP)
The concept of utilizing segmentation masks for pixel pooling can be integrated into various LDGD frameworks that jointly predict segmentation masks and grasp maps. We illustrate this integration using the CROG framework [46], a robust LDGD model that leverages CLIP-based visual and textual encoders combined through cross-attention mechanisms to fuse vision-language features. CROG employs a multi-modal Feature Pyramid Network (FPN) followed by additional cross-attention layers that incorporate textual embeddings. These enriched features are subsequently passed through a multi-task projector, which simultaneously outputs binary segmentation masks and grasp parameter maps. The projector architecture consists of sequential Conv2D and upsampling layers designed to maintain feature dimensions consistent with the original input images.
We introduce Mask-Guided Feature Pooling (MFP) within the multi-task projector (see Fig. 2). Our training methodology comprises two sequential stages: i) Pretraining the Segmentation Mask Branch: In this initial stage, we freeze the grasp prediction heads and train the segmentation mask branch separately under a standard Referring Expression Segmentation (RES) training setup. This stage intentionally allows slight overfitting to ensure high-quality object-level masks, facilitating precise pixel pooling in subsequent training. ii) Mask-Guided Grasp Prediction: In this stage, we introduce mask-guided pooling connections within the multi-task projector, focusing the grasp predictions exclusively on visually grounded object regions defined by the segmentation masks. These pooled features undergo further refinement via a simple Multi-Layer Perceptron (MLP) to generate accurate grasp maps.
Two critical design choices ensure robustness. First, we do not use ground-truth object masks for pooling, as this would make grasp predictions overtly sensitive to mask accuracy, which is challenging in cluttered scenes with complex referring text. Second, we avoid using open-set detectors during training, since state-of-the-art object mask segmentation models currently achieve only 50–60% IoU on our benchmarks, which is insufficient for reliable pooling.
3.5 Loss Function
Prior work utilizes the Smooth L1 loss for pixel-wise prediction in visual grounding, penalizing smaller errors quadratically and larger errors linearly, defined per pixel $(\hat{x}_{p},x_{p})$ as:
$L_{\mathrm{smoothL1}}(\hat{x}_{p},x_{p})=\begin{cases}\tfrac{(\hat{x}_{p}-x_{p})^{2}}{2\beta},&\!\text{if }|\hat{x}_{p}-x_{p}|<\beta,\\ |\hat{x}_{p}-x_{p}|-\tfrac{\beta}{2},&\!\text{otherwise.}\end{cases}$ (3)
We extend this to formulate weighted Smooth L1 loss $w_{p}\,L_{\mathrm{smoothL1}}(\hat{x}_{p},x_{p})$, by assigning per‐pixel weights $w_{p}=1+\alpha\,Q_{\mathrm{gt},p}$, where $Q_{\mathrm{gt},p}\in[0,1]$ is the ground‐truth quality map indicating the object grasp region. By emphasizing suitable grasp regions, we enhance precise grasp estimation. Our experiments demonstrate that this modified loss function is beneficial in low-resource settings, where the dataset contains only single grasp annotation per object, ensuring more reliable and efficient learning.
3.6 Inverse Kinematics Control for Manipulation
Our approach for real world object manipulation transforms the 2D grasp prediction into a desired 3D pose. Specifically, we obtain the final grasp pose in the robot’s base frame, $\mathbf{T}_{\text{base}}^{\text{pose}}$, by composing three transformations:
$\mathbf{T}_{\text{base}}^{\text{pose}}\;=\;\mathbf{T}_{\text{base}}^{\text{ee}}\;\mathbf{T}_{\text{ee}}^{\text{cam}}\;\mathbf{T}_{\text{cam}}^{\text{pose}},$ (4)
where:
  1. 1.
    $\mathbf{T}_{\text{base}}^{\text{ee}}$ is the forward-kinematics transformation from the robot base to its end-effector,
  2. 2.
    $\mathbf{T}_{\text{ee}}^{\text{cam}}$ is the hand–eye calibration matrix (mapping from end-effector to camera),
  3. 3.
    $\mathbf{T}_{\text{cam}}^{\text{pose}}$ is the camera-to-pose transformation predicted by our grasp model.
To execute a grasp in closed loop, we generate a smooth interpolation $\mathcal{C}(t)$ in $\mathrm{SE}(3)$ between the initial end-effector pose $\mathbf{T}_{\text{init}}$ and the desired grasp pose $\mathbf{T}_{\text{goal}}$, as:
$\mathcal{C}(t)\;:\;[0,\,1]\;\to\;\mathrm{SE}(3),\quad\mathcal{C}(0)=\mathbf{T}_{\text{init}},\quad\mathcal{C}(1)=\mathbf{T}_{\text{goal}}.$ (5)
At each control step $t$, the desired pose is
$\mathbf{T}_{\text{des}}\;=\;\mathcal{C}(t).$ (6)
Let $\mathbf{T}_{\text{current}}$ denote the current pose of the end effector. We define the positional error
$\mathbf{e}_{p}=\mathbf{p}_{\text{des}}-\mathbf{p}_{\text{current}},$ (7)
and compute the rotational error using the matrix logarithm of the orientation difference:
$\mathbf{R}_{\text{err}}=\mathbf{R}_{\text{des}}\,\mathbf{R}_{\text{current}}^{\top},\qquad\boldsymbol{\theta}_{\text{err}}=\mathrm{Log}\!\bigl{(}\mathbf{R}_{\text{err}}\bigr{)},$ (8)
where $\mathbf{R}_{\text{des}}$ and $\mathbf{R}_{\text{current}}$ are the rotation components of $\mathbf{T}_{\text{des}}$ and $\mathbf{T}_{\text{current}}$, respectively, and $\mathrm{Log}(\cdot)$ is the matrix logarithm on $\mathrm{SO}(3)$. We form the combined error vector
$\mathbf{e}=\begin{bmatrix}\mathbf{e}_{p}\\ \boldsymbol{\theta}_{\text{err}}\end{bmatrix}\;\in\;\mathbb{R}^{6}.$ (9)
The desired end-effector velocity is then
$\dot{\mathbf{x}}_{\text{target}}\;=\;K_{p}\;\mathbf{e},$ (10)
where $K_{p}$ is a diagonal gain matrix. We compute the joint velocity command by projecting $\dot{\mathbf{x}}_{\text{target}}$ through the robot Jacobian pseudoinverse:
$\dot{\mathbf{q}}_{\text{cmd}}\;=\;\mathbf{J}^{\dagger}\;\dot{\mathbf{x}}_{\text{target}}.$ (11)
Finally, we update the robot’s joint states with $\dot{\mathbf{q}}_{\text{cmd}}$ and advance $t\!\leftarrow\!t+\Delta t$. This process repeats until
$\|\mathbf{T}_{\text{des}}-\mathbf{T}_{\text{current}}\|\;\leq\;\epsilon,$ (12)
indicating successful convergence to the target grasp pose.
4 RefGraspNet Dataset: Learning to Grasp from 200M+ Grasp Poses
Dataset Real/Simulated No. of Images Ref. Text Unique Obj. Grasp Poses Obj. Masks Avg. Obj./Img.
Jacquard [8] 54K 967K 1
VMRD [60] 4.5K 31 51.5K 3.5
OV-Grasp [23] Both 63K 117
ACRONYM [10] 262 17.7M
RoboRefIt [32] 10.7K 50.7K 66 7.1
OCID-VLG [46] 1.7K 89.6K 31 521K 17.08
GraspAnything++ [49] 994K 10M 236 33M 3.4
Ours: RefGraspNet 97K 12.25M 88 219M 9.57
Table 1: Comparison of existing 4-DoF LDRG datasets. RefGraspNet provides 200M+ high quality grasps, object masks for each scene, and thus can be used to finetune both referring segmentation and grasping models.
We build on existing research and datasets in language driven grasping by creating a new corpus of real world cluttered images and their grasp annotations. Our dataset, RefGraspNet, is derived from the GraspNet-1B dataset [11] and comprises a diverse corpus collected using 88 unique objects arranged in cluttered scenes. We construct 10 instruction templates such as “Grasp the {object}”, “Find the {object} in this scene”, etc. and expand the object names to create referring expressions. In scenes with multiple objects, we leverage an open-source model, DeepSeek-VL [30], to provide additional attributes such as color, shape and relative position. We chose GraspNet-1B as our base corpus since it consists of 1 billion + grasp poses annotated with sampling and force-closure techniques. We sample grasps with a confidence score greater than 70%, and our final corpus comprises 219M grasp poses, 8 times bigger than the previous largest dataset for language driven 4 DoF grasping. Comparisons with prior datasets are provided in Tab. 1. We randomly choose 70% of the objects to be seen objects (with the rest as unseen) and create train/val/test-seen and test-unseen splits for our experiments.
5 Experiments
Here, we present our experiments evaluating MapleGrasp on OCID-VLG and RefGraspNet benchmarks, while also comparing our framework to prior LDRG and VLA methods in simulation and real-world setups.
5.1 OCID-VLG Benchmark
The OCID-VLG corpus [46] contains 1763 highly cluttered indoor tabletop RGB-D scenes with 31 unique objects. Objects have annotated 4-DoF grasp poses and segmentation masks, associated with referring expressions describing attributes such as color, shape, and relative position. With 89K unique (RGB-Text-Mask) tuples, it is the largest publicly available benchmark for language-driven grasp detection frameworks [^1] GraspAnything++ is another open-source corpus, but the publicly shared subset is significantly smaller than what was used for training their models, preventing fair benchmarking.. We evaluate various CLIP-based architectures on OCID-VLG, applying the MapleGrasp framework to Ref2Grab [2], which uses frozen CLIP encoders for improved generalization, and CROG [46], as detailed in Sec. 3.4. Results are summarized in Tab. 2.
Model IOU Pr@50 Pr@60 Pr@70 Pr@80 Pr90 J@1 J@Any
DetSeg + CLIP 29 27.2 20.9 17.5 17.2 16 28.1 39.2
GR-ConvNet + CLIP 31.3 21.0 11.6 5.5 2.4 0.5 9.73 15.41
SSG + CLIP 33.6 35.6 35.6 35.5 35.5 32.8 33.5 34.7
CROG 81.1 96.9 94.8 87.2 64.1 16.4 77.2 87.7
Ref2Grab 83.26 85.68 85.13 91.53 89.69 83.21 70.5 79.1
ETRG 80.11 - - - - - 82.28 91.12
MapleGrasp-Ref2Grab 83.78 85.96 85.41 91.74 90.13 83.79 76.8 84.7
MapleGrasp-CROG 81.36 97.40 95.32 87.9 65.4 16.4 86.15 91.9
Table 2: Results on OCID-VLG. We implement all baseline methods except ETRG, whose results are taken from their paper. Our comparison is limited to methods that do not utilize depth maps during training.
Referring Text Molmo+SAM+GraspNet CROG MapleGrasp
Grasp the rightmost red food box [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Grab the transparent food bag [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Find the banana furthest away from me [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 3: Qualitative comparisons of LDRG approaches under complex referring queries. The zero-shot baseline (first column) struggles to identify referred object due to distractors. MapleGrasp (last column) can predict a stable and accessible grasp for each queried object.
Applying mask proposals enhances both the CROG and Ref2Grab baselines in terms of the Jacquard Index J@N metric, which evaluates the top-N grasp rectangle candidates that exhibit a rotation angle difference of less than $30^{\circ}$ and an IoU greater than 0.25 compared to the ground-truth grasp rectangle. MapleGrasp achieves a 12% relative improvement in J@1 over previous state-of-the-art methods, demonstrating higher precision in grasp prediction. However, referring expression segmentation (RES) accuracy, measured using Prec@X, which quantifies the percentage of test cases where the IoU exceeds $X\in\{0.5,0.6,0.7,0.8,0.9\}$, remains relatively unchanged. This may be due to CROG losing pretrained knowledge during CLIP layer fine-tuning. In contrast, MapleGrasp-Ref2Grab, which employs frozen vision-language encoders, achieves higher performance in the RES task but struggles with grasp map prediction due to its limited trainable parameter space (4M vs. CROG’s 167M). Overall, our findings suggest that MFP enhances both frozen and fully fine-tuned CLIP-based models for language-driven grasp detection.
5.2 Ablation Studies
We conduct ablation studies on MapleGrasp (LABEL:tab:ablations) to analyze its effects due to architectural modifications. Results highlight the importance of accurate RES for mask pooling; using Molmo, which achieves lower IoU on RES tasks, negatively impacts grasp accuracy due to complex referring queries requiring task-specific visual grounding. While using frozen CLIP improves RES, it decreases grasp performance because of fewer trainable parameters. Replacing cross-attention with MLP-Mixers, despite reported benefits [44], does not significantly improve performance. Furthermore, replacing the weighted smooth L1 loss with a standard Smooth L1 loss leads to reduced scores, likely from slower convergence. All ablations use a fixed number of training epochs for consistency.
Method IoU J@1 J@Any
MapleGrasp-CROG 81.36 86.15 91.90
(1) Molmo+SAM for Mask Pooling 73.42 77.79 80.16
(2) Frozen CLIP Layers 83.52 72.19 74.44
(3) No Cross-Attn, MLP-Mixer 80.16 84.37 90.45
(4) Standard Smooth L1 Loss 80.17 83.23 89.18
5.3 RefGraspNet Benchmark
We further evaluate previous baselines and MFP-enhanced architectures on our RefGraspNet dataset, establishing a strong benchmark for future research. The dataset comprises 88 unique objects, with 70% assigned as seen objects and 30% as unseen objects. Results for both test splits are reported in Tab. 4. Similar trends as observed in the OCID-VLG benchmark hold here, with MFP providing significant performance gains for both Ref2Grab and CROG. However, all trained models perform worse on unseen objects, highlighting a lack of zero-shot generalization. In contrast, open-set object detectors such as Molmo [7] achieve higher test unseen IoU. This motivates us to incorporate Molmo in an ensembling setup to boost zero-shot segmentation, which is essential for robust grasp pose estimation.
Test Seen Test Unseen
Model IoU J@1 J@Any IoU J@1 J@Any
DetSeg + CLIP 48.18 40.19 41.56 30.57 27.18 27.20
GR-ConvNet +
CLIP 52.34 44.19 46.56 32.18 30.19 32.17
SSG + CLIP 57.16 48.78 50.10 38.19 34.16 34.89
CROG 84.49 85.32 86.49 69.95 70.81 71.99
Ref2Grab 87.32 73.27 75.45 75.67 62.12 63.28
MapleGrasp-Ref2Grab 87.74 75.67 77.89 75.11 65.19 67.22
MapleGrasp-CROG 84.86 89.15 89.67 70.43 74.92 75.67
Table 4: Results on RefGraspNet. The MFP framework leads to consistent improves across seen and unseen objects. However, performance on unseen objects deteriorates for all methods.
5.4 Cross-Dataset Testing
To assess the generalizability of models trained on RefGraspNet, we cross-test them on the OCID-VLG benchmark. We compare MapleGrasp-Ref2Grab and MapleGrasp-CROG, and present results in Tab. 5. Our findings reveal that models trained on RefGraspNet generalize better than those trained solely on OCID-VLG. The dataset’s diversity in grasp poses and text expressions enhances inference on other benchmarks. This highlights the advantages of training on RefGraspNet, encouraging further adoption of our corpus to push performance boundaries.
OCID-VLG RefGraspNet
Model Train-Dataset J@1 J@Any J@1 J@Any
MapleGrasp-Ref2Gr OCID-VLG 76.8 84.7 41.8 42.9
RefGraspNet 71.2 74.4 75.7 77.9
MapleGrasp-CROG OCID-VLG 86.2 91.9 33.2 35.6
RefGraspNet 68.2 73.8 89.2 89.7
Table 5: Cross-dataset evaluation results. Models trained on RefGraspNet generalize better than OCID-VLG.
5.5 LIBERO Simulation Experiments
We deploy the MapleGrasp model trained on RefGraspNet and extend grasp poses to 3D using an inverse kinematics controller for trajectory prediction (Sec. 3.6). This setup allows direct comparisons with Vision-Language-Action (VLA) models. Since our approach focuses on language-driven grasp pose estimation, we cannot evaluate task success rates directly. Instead, we define grasp success as trials in which an object is lifted and held above ground level for at least two seconds. We report J@1 and J@Any metrics. Experiments are conducted on three task suites from the LIBERO benchmark: SPATIAL, GOAL, and OBJECT. Since VLAs are fine-tuned on 50 teleoperated demonstrations per task, we automate 4-DoF grasp annotations using forward kinematics-based 3D grasp pose estimation and project the masks to 2D, inverting steps described in Sec. 3.1. We fine-tuned MapleGrasp on these annotations to adapt to the simulation environment. At inference time, MapleGrasp predicts 2D grasps, which are transformed and executed in 3D by the controller. Results in Tab. 6 compare our method with VLA and diffusion baselines.
Model SPATIAL GOAL OBJECT
Diffusion Policy 0.77 0.69 0.92
Octo 0.78 0.84 0.88
OpenVLA 0.84 0.79 0.90
MapleGrasp-CROG 0.87 0.85 0.90
Table 6: Libero simulation results across SPATIAL, GOAL and OBJECT task suites. Scores are reported using J@1.
Our results demonstrate that our grasping+control pipeline achieves similar accuracy to recent VLAs in grasping tasks. Interestingly, decoupling grasping and control improves generalization, as trajectory-based imitation learning methods fail to complete unseen tasks. However, our model successfully determines grasp poses for similar but unseen objects, allowing the control algorithm to generate a stable trajectory without requiring video demonstrations. Cross-dataset testing in Tab. 7 reveals the capabilities of our approach in handling unseen queries across task suites. Since our LDRG model is only concerned with predicting grasp poses and we leverage a controller to determine the trajectory, our method can be deployed without learning from tele-operated demonstrations.
Model Train-Dataset GOAL OBJECT
OpenVLA GOAL 0.79 0.00
OBJECT 0.00 0.90
MapleGrasp-CROG GOAL 0.85 0.62
OBJECT 0.68 0.90
Table 7: Comparison of OpenVLA and our method in cross-dataset settings. OpenVLA struggles with seen objects under novel tasks, while our method generalizes well to unseen queries.
5.6 Real-Robot Experiments
Figure 3: Setup for robot trials using 7 DoF franka arm and our modular pipeline described in section Sec. 3.6
We evaluate various LDRG methods in a real-world robotic setup using a Franka arm with a mounted camera at the gripper, as shown in Fig. 3. We construct diverse tabletop scenes featuring single and multiple objects, distractors, and varying lighting conditions. Results are presented in Tab. 8. For real-robot trials, MapleGrasp integrates Molmo-based object mask refinement, which compares our model’s segmentation predictions with Molmo’s out-of-vocabulary predictions, selecting the highest mutual overlap. MapleGrasp outperforms other baselines, achieving a notable 7% relative improvement in cluttered setups.
Model Single Cluttered
Det-Seg+CLIP 0.27 0.21
GR-ConvNet+CLIP 0.30 0.27
Molmo+SAM+GraspNet 0.45 0.47
SSG+CLIP 0.21 0.17
Ref2Grab 0.42 0.40
CROG 0.55 0.53
MapleGrasp-CROG 0.59 0.57
Table 8: Real robot experiments. Success rate measured across 10 trials for every grasp instruction.
6 Conclusion
Our results highlight mask-guided feature pooling as a simple yet effective technique for language-driven robotic grasping. By selectively pooling predictions within target regions, MapleGrasp reduces misidentifications and grasp errors in cluttered environments. Our two-stage training strategy, which first predicts segmentation masks and subsequently refines grasp predictions, further enhances accuracy and computational efficiency. Compared to end-to-end vision-language-action models, our modular approach demonstrates superior scalability and robustness, and enables rapid adaptation using readily available vision-language models. Moreover, these systems eliminate the need for time-intensive teleoperated demonstrations and can be customized using a small corpus of RGB-text-grasp annotations. Real-world validation confirms significant improvements in grasp accuracy and efficiency.

### v2

@@
 Foundational models, particularly Vision-Language Models (VLMs), aim to bridge this gap by leveraging extensive pre-trained multi-modal representations to enhance robotic perception and interaction [23, 62].
-An end-to-end robotic manipulation framework typically comprises four key stages: task planning, visual grounding, object manipulation, and goal-oriented placement.
+An autonomous robotic manipulation framework typically comprises four key stages: task planning, visual grounding, object manipulation, and goal-oriented placement.
 Recent advances have demonstrated the effectiveness of VLMs across these stages, incorporating large-scale pre-trained representations to improve generalization [51].
@@
 1.
-While VLAs have shown remarkable performance in long-range tasks across both simulation and real-world settings, their deployment in novel environments presents three key challenges: i) Task-Specific Data Collection: Collecting teleoperated or human demonstrations is often difficult and time-consuming for novel interactions [53]; ii) Expensive fine-tuning: Large-scale VLAs require task-specific fine-tuning, which is computationally expensive.
-Although parameter-efficient fine-tuning (PEFT) techniques alleviate some costs, they restrict adaptability by modifying only a small subset of parameters [20]; iii) Limited generalization: Fine-tuned VLAs often fail to generalize across dynamic environments, particularly when objects are repositioned or when visually similar distractor objects are introduced [46].
+While VLAs have shown remarkable performance in tasks across both simulation and real-world settings, their deployment in novel environments presents three key challenges: i) Task-Specific Data Collection: Collecting teleoperated or human demonstrations is often difficult and time-consuming for novel interactions [50]; ii) Expensive fine-tuning: Large VLAs require task-specific fine-tuning, which is computationally expensive requiring hundreds of GPU hours.
+Although parameter-efficient fine-tuning (PEFT) techniques alleviate some costs, they restrict adaptability by modifying only a small subset of parameters [18]; iii) Limited generalization: Fine-tuned VLAs often fail to generalize in dynamic environments, particularly when objects are repositioned or when visually similar distractor objects are introduced [44].
 Given these challenges, modular frameworks offer an attractive alternative that balances generalizability with ease of deployment.
 By decoupling perception, planning, and manipulation, modular architectures allow state-of-the-art models in each component to be independently improved and integrated as research advances.
-Furthermore, LLM-agentic frameworks can serve as high-level decision-makers, dynamically invoking perception and manipulation modules for task execution in novel environments [15, 41, 42].
+Furthermore, LLM-agentic frameworks can serve as high-level decision-makers, dynamically invoked perception and manipulation modules for task execution in novel environments [13, 36, 38].
 Compounded errors within modular frameworks can be mitigated using closed-loop feedback and error planning [5, 14].
 In this work, we focus on a critical capability within modular frameworks, language-driven grasp detection.
-This task involves predicting a stable grasp pose for a specified object based on RGB-D images and natural language referring expressions.
-We focus on 2D grasp detection, which can be naturally extended to 3D using depth information and camera intrinsic parameters.
+This task predicts a stable grasp pose for a specified object based on RGB-D images and a free-form language expression.
+An accurate LDRG model can enable low-level execution for high-level LLM planners, advancing fully autonomous agentic frameworks for human-robot interaction.
+Our proposed method, MapleGrasp, can predict both 4 DoF top-down and oriented 6 DoF grasps.
 Our contributions include:
   1. 1.
-    A novel two-stage training strategy for LDRG: initially predicting text-referred object masks, then refining grasp predictions via mask-guided feature pooling.
+    A novel two-stage training architecture for LDRG: initial text-referred object mask prediction, followed by mask-guided feature pooling for grasp refinement.
   2. 2.
@@
   3. 3.
-    Introduce RefGraspNet, a large-scale open-source dataset comprising over 200 million grasp annotations, eight times larger than existing datasets, consisting of challenging real-world scenarios for grasping.
+    Introduce RefGraspNet, a large-scale open-source benchmark comprising over 200 million grasps, eight times larger than existing datasets (Tab.
+1), consisting of challenging real-world scenarios for grasping.
   4. 4.
-    Extend 2D grasp predictions to 3D via depth maps and inverse kinematics, achieving performance comparable to recent Vision Language Action (VLA) models on the LIBERO benchmark, with superior generalization to diverse tasks while being parameter efficient.
+[…]
 2 Related Works
@@
 The 6-DoF representation extends this to 3D, incorporating three degrees of translation and three degrees of rotation required for stable object grasping [29, 9, 39].
-Early approaches to object grasping relied on kinematics and geometric methods to identify stable grasp poses from depth maps [3] and 3D object models, using datasets generated through offline labeled training data, human demonstrations, or trial-and-error methods [13, 9].
 The use of RGB-D data has shown promising results for precision grasping, where CNN and ResNet-based architectures leverage visual features to predict grasp pose parameters [20, 46, 63, 56].
-Additionally, 3D-CNN models have been employed to extract 3D features from point clouds [39, 38, 51].
 Recent advancements in transformer architectures and self-supervised learning have bridged the gap between vision and textual modalities, shifting research towards incorporating language as an additional modality for object manipulation [51].
 Language-driven grasp detection (LDGD) refers to the task of associating referring text expressions with RGB-D images to predict a suitable robotic grasp pose for object manipulation [47].
-Two primary approaches have emerged for LDGD: i) Visual grounding of referring expressions: This approach predicts segmentation masks for the target object [7, 55, 61], followed by the application of pre-trained grasp detection networks to compute grasp poses [2, 27, 32].
-ii) End-to-end grasp pose estimation: This method directly predicts grasp poses from input text and RGB-D modalities [56, 63, 52].
-Figure 2: Our framework trains on fused vision-language embeddings for object segmentation, then applies mask pooling in grasp prediction heads.
-MapleGrasp is an network-agnostic MFP block with no extra parameters and can be used with the CROG backbone as shown.
+Two primary approaches have emerged for LDGD: i) Visual grounding of referring expressions: Predicts segmentation masks for the target object [6, 53, 60], followed by pre-trained grasp detection networks to compute grasp poses [2, 25, 30].
+ii) End-to-end grasp pose estimation: Directly predicts grasp poses from input text and RGB-D images [54, 61, 49].
+Figure 2: MapleGrasp: Our framework trains on fused vision-language embeddings for object segmentation, then applies mask pooling in grasp prediction heads.
+The predicted grasp maps can be used for both top-down and oriented grasping.
 End-to-End Language-Driven Grasping.
-Our work focuses on utilizing RGB images of a scene along with free-form referring expressions to determine a stable 4-DoF grasp for the target object in cluttered environments.
-We extend this by incorporating depth maps to transform the 2D grasp into 3D and leveraging a differential inverse kinematics controller to compute the joint-space trajectory for grasp execution.
-Several studies have explored similar setups for LDGD [45, 46].
 Recent work has introduced diffusion models that learn a probabilistic diffusion process to generate grasp poses during inference [47, 32, 33].
 Vision-language fusion provides an alternative approach, offering faster convergence and lower inference times [31, 22, 51].
-The ETRG framework enhances vision-language feature fusion for grasp detection by integrating vision-language adapters into frozen CLIP vision and text encoders [57].
-MaskGrasp employs a tri-head architecture with cross-attention mechanisms using text embeddings, region-of-interest (ROI) feature embeddings, and mask embeddings [47].
+The ETRG framework enhances vision-language feature fusion for grasp detection by integrating trainable adapters into frozen CLIP vision and text encoders [55].
+MaskGrasp employs a tri-head architecture with cross-attention mechanisms using text embeddings, region-of-interest feature embeddings, and mask embeddings [45].
 GraspSAM adapts the Segment Anything Model [19] for grasp detection and utilizes GroundingDINO [26] to transform text expressions into object bounding boxes [34].
-In our work, we analyze the impact of mask pooling on grasp detection and demonstrate that combining mask pooling with open-set object detectors significantly enhances performance on unseen objects—an issue that previous studies have not adequately addressed.
-Moreover, we present the first exhaustive comparison of LDRG approaches against large Vision-Language-Action (VLA) models for language-guided grasping tasks.
+GraspMolmo [8] trains the Molmo VLM on the grasp point detection task followed by a grasp proposal sampler to identify accurate grasps for task oriented grasping.
+Our proposed architecture predicts grasp pose using quality, width and angle maps.
+Several studies have explored similar setups for LDGD [43, 44, 41].
+These maps can be used to predict both 4 DoF and 6 DoF grasps by computing the grasp point on the object and overlapping with an independent grasp proposal network.
+In this work, we study mask pooling for grasp detection and demonstrate that localizing the grasp prior to predicting its angle and width leads to a more efficient learning process.
+This approach improves both training efficiency and test-time accuracy.
+Additionally, we present the first comprehensive comparison of LDRG approaches with large VLAs in the context of language-guided grasping tasks.
 Vision-Language-Action Models.
@@
 However, the large model sizes of VLAs often demand substantial GPU resources for fine-tuning and real-world deployment [48].
-Recent research has sought to mitigate these computational challenges through parameter-efficient training techniques and action chunking methods [19, 20], enabling deployment on smaller GPUs with real-time kernel compatibility for seamless robotic control.
-3 Mask-Guided Feature Pooling for Language Driven Grasp Detection
-In this section, we introduce our approach of leveraging segmentation masks for feature pooling to focus on relevant regions in an image for targeted grasp pose estimation in cluttered environments.
-Furthermore, we present our two-stage training framework and discuss our improved training objective used to optimize our method.
+Recent research has sought to mitigate these computational challenges through parameter-efficient training techniques and action chunking [17, 18], enabling deployment on smaller GPUs with real-time kernel compatibility.
+3 Methodology
+We propose object-mask feature pooling to isolate relevant regions for grasp detection in clutter, detail our architecture (Fig. 2), and outline an enhanced training objective.
 3.1 Language-Driven Grasp Detection
-Given an RGB-D image and a corresponding referring expression, language-driven grasping models predict a stable grasp pose to pick up the specified object using either a parallel gripper or a dexterous robotic hand.
-This task poses three primary challenges: i) accurately identifying the target object using textual cues provided by the referring expression, such as color, shape, and relative positioning; ii) determining a stable and collision-free grasp pose that can be reliably executed in real-world environments without interference from surrounding objects; and iii) generalizing grasp predictions effectively to previously unseen objects that were not included in the training dataset.
-We focus on grasp pose estimation in 2D, where where grasps are typically represented as rectangles [49, 37, 34].
-Each 2D grasp is defined by the center coordinates $(x,y)$ of the grasp rectangle, the grasp width $w$, and the orientation angle $\theta$ relative to a fixed horizontal axis.
-Given that this study focuses on top-down grasps performed with a parallel gripper, the height of the grasp rectangle remains constant.
-Such 2D grasp representations naturally extend into 3D space by incorporating depth information.
-Formally, a 2D grasp $\mathbf{g_{2D}}=(x,\,y,\,w,\,\theta)$ is converted into a 3D grasp pose $\mathbf{T_{3D}}$ by projecting the pixel $(x,y)$ into 3D space.
-With depth value $D(x,y)$ at pixel $(x,y)$, the 3D coordinates relative to the camera are computed as:
-$Z=D(x,y),\quad X=\frac{(x-c_{x})\,Z}{f_{x}},\quad Y=\frac{(y-c_{y})\,Z}{f_{y}},$ (1)
-where $f_{x},f_{y}$ denote the camera focal lengths, and $c_{x},c_{y}$ represent the principal point offsets.
-The computed coordinates $(X,Y,Z)$ correspond to the grasp center position in the camera frame.
-The resulting top-down 3D grasp pose transformation matrix $\mathbf{T}_{3D}$ is then defined as:
-$\mathbf{T}_{3D}=\begin{bmatrix}\cos\theta&-\sin\theta&0&X\\ \sin\theta&\cos\theta&0&Y\\ 0&0&1&Z\\ 0&0&0&1\end{bmatrix}.$ (2)
-Direct prediction of all four grasp parameters (location, width, and orientation) at every pixel often yields suboptimal results due to their weak direct correlation with the original RGB-D data.
-Thus, contemporary approaches typically estimate grasp parameters using three distinct 2D prediction maps: i) Grasp Quality Map: Assigns a pixel-wise score for grasp quality; the pixel with the highest score represents the best grasp location, ii) Grasp Width Map: Encodes the optimal grasp width at each pixel and iii) Grasp Angle Map: Stores the grasp orientation $\theta$ at each pixel.
-This decoupled representation enhances the reliability and interpretability of grasp prediction models.
-3.2 Referring Expression Segmentation
-Language-Driven Grasp Detection (LDGD) shares strong similarities with a well-established computer vision task known as Referring Expression Segmentation (RES).
-The goal of RES is to predict a binary segmentation mask that separates the referred object from its surroundings [33, 62].
-Since both tasks involve object-level visual grounding, several LDGD frameworks incorporate both segmentation mask prediction and grasp rectangle estimation from the referring query [47, 57].
-These architectures typically contain separate projection layers for predicting the binary object mask and the grasp maps.
-However, since predictions are made independently, the segmentation mask may be correctly identified while grasp maps remain inaccurate.
+From an RGB-D image and a referring expression identifying a target object to grasp, language-driven grasping models must predict a stable grasp pose for a parallel gripper or dexterous hand.
+This poses three primary challenges: (i) identifying the target from textual cues (color, shape, spatial relations); (ii) selecting a collision-free grasp executable in real settings; and (iii) generalizing to unseen objects.
+Vision-language driven grasping is studied with both 4 DoF top-down and 6 DoF oriented grasps.
+While 6 DoF grasps are more accurate, annotating such grasp poses at a large scale to create real-world datasets is challenging.
+A number of works thus rely on image-based datasets with 4 DoF grasp rectangles.
+We design our method such that it can be used with both 4 DoF and 6 DoF grasping datasets.
+Fig.
+3 illustrates the two grasp formats and their transformation via an external proposal sampler.
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 3: Conversion from a 4 DoF to a 6 DoF grasp pose: (top-left) initial top-down grasp rectangle, (top-right) sampled contact points reprojected to 2D, (bottom-left) 6 DoF grasp proposals from Contact-GraspNet [40], (bottom-right) selected 6 DoF grasp.
+Formally, MapleGrasp predicts three dense maps, each of size $H\times W$: a grasp quality map $Q\in\mathbb{R}^{H\times W}$, a grasp width map $w\in\mathbb{R}^{H\times W}$, and a grasp angle map $\theta\in\mathbb{R}^{H\times W}$.
+The quality map assigns a score to each pixel, indicating the likelihood of a successful grasp.
+The width map encodes the optimal gripper width at each pixel, while the angle map represents the in-plane grasp orientation $\theta\in[-\pi/2,\pi/2]$.
+A 2D rectangular grasp $g_{\text{2D}}=(x^{*},y^{*},w^{*},\theta^{*})$ is extracted by selecting the pixel $(x^{*},y^{*})=\arg\max_{(i,j)}Q_{i,j}$, and reading $w^{*}=w_{x^{*},y^{*}}$, $\theta^{*}=\theta_{x^{*},y^{*}}$.
+This rectangular grasp is projected onto the point cloud $\mathcal{P}$ to identify a subset of points $\mathcal{P}_{g}\subset\mathcal{P}$ within the grasp rectangle.
+A set of candidate 6-DoF grasp poses $\{G_{k}\}$ is generated using an independent grasp sampler.
+Each candidate pose $G_{k}$ is evaluated by computing its intersection with $\mathcal{P}_{g}$, and the final grasp $G^{*}$ is chosen as the one that overlaps the maximum number of graspable points: $G^{*}=\arg\max_{k}\,|\mathcal{P}_{g}\cap\mathcal{V}(G_{k})|$, where $\mathcal{V}(G_{k})$ denotes volume swept by the gripper pose $G_{k}$.
 3.2 Motivation for Mask Pooling
-The predicted binary segmentation masks highlight the image regions corresponding to the referring expression, providing an implicit constraint for grasp pose estimation.
-If the segmentation mask is correctly predicted, then the optimal grasp pose should lie within the segmented region.
-This suggests that the grasp prediction module does not need to consider grasp locations outside the segmented object, provided that the segmentation branch can accurately localize the object.
-With recent advancements in VLMs for RES, zero-shot object segmentation has significantly improved.
-Consequently, off-the-shelf segmentation models can be used to extract masks for grasp pose estimation networks to predict stable grasp poses [7, 61].
-We hypothesize that incorporating segmentation masks as a guiding mechanism within the grasp detection pipeline can improve accuracy by focusing predictions on relevant regions.
-Prior work has not explored this direction, and we argue that integrating mask pooling with LDRG can effectively utilize advances in RES to enhance grasp detection accuracy.
-3.4 Mask-Guided Feature Pooling (MFP)
-The concept of utilizing segmentation masks for pixel pooling can be integrated into various LDGD frameworks that jointly predict segmentation masks and grasp maps.
-We illustrate this integration using the CROG framework [46], a robust LDGD model that leverages CLIP-based visual and textual encoders combined through cross-attention mechanisms to fuse vision-language features.
-CROG employs a multi-modal Feature Pyramid Network (FPN) followed by additional cross-attention layers that incorporate textual embeddings.
-These enriched features are subsequently passed through a multi-task projector, which simultaneously outputs binary segmentation masks and grasp parameter maps.
-The projector architecture consists of sequential Conv2D and upsampling layers designed to maintain feature dimensions consistent with the original input images.
-We introduce Mask-Guided Feature Pooling (MFP) within the multi-task projector (see Fig.
-2).
-Our training methodology comprises two sequential stages: i) Pretraining the Segmentation Mask Branch: In this initial stage, we freeze the grasp prediction heads and train the segmentation mask branch separately under a standard Referring Expression Segmentation (RES) training setup.
-This stage intentionally allows slight overfitting to ensure high-quality object-level masks, facilitating precise pixel pooling in subsequent training.
-ii) Mask-Guided Grasp Prediction: In this stage, we introduce mask-guided pooling connections within the multi-task projector, focusing the grasp predictions exclusively on visually grounded object regions defined by the segmentation masks.
-These pooled features undergo further refinement via a simple Multi-Layer Perceptron (MLP) to generate accurate grasp maps.
-Two critical design choices ensure robustness.
-First, we do not use ground-truth object masks for pooling, as this would make grasp predictions overtly sensitive to mask accuracy, which is challenging in cluttered scenes with complex referring text.
-Second, we avoid using open-set detectors during training, since state-of-the-art object mask segmentation models currently achieve only 50–60% IoU on our benchmarks, which is insufficient for reliable pooling.
-3.5 Loss Function
-Prior work utilizes the Smooth L1 loss for pixel-wise prediction in visual grounding, penalizing smaller errors quadratically and larger errors linearly, defined per pixel $(\hat{x}_{p},x_{p})$ as:
-$L_{\mathrm{smoothL1}}(\hat{x}_{p},x_{p})=\begin{cases}\tfrac{(\hat{x}_{p}-x_{p})^{2}}{2\beta},&\!\text{if }|\hat{x}_{p}-x_{p}|<\beta,\\ |\hat{x}_{p}-x_{p}|-\tfrac{\beta}{2},&\!\text{otherwise.}\end{cases}$ (3)
-We extend this to formulate weighted Smooth L1 loss $w_{p}\,L_{\mathrm{smoothL1}}(\hat{x}_{p},x_{p})$, by assigning per‐pixel weights $w_{p}=1+\alpha\,Q_{\mathrm{gt},p}$, where $Q_{\mathrm{gt},p}\in[0,1]$ is the ground‐truth quality map indicating the object grasp region.
-By emphasizing suitable grasp regions, we enhance precise grasp estimation.
-Our experiments demonstrate that this modified loss function is beneficial in low-resource settings, where the dataset contains only single grasp annotation per object, ensuring more reliable and efficient learning.
-3.6 Inverse Kinematics Control for Manipulation
-Our approach for real world object manipulation transforms the 2D grasp prediction into a desired 3D pose.
-Specifically, we obtain the final grasp pose in the robot’s base frame, $\mathbf{T}_{\text{base}}^{\text{pose}}$, by composing three transformations:
-$\mathbf{T}_{\text{base}}^{\text{pose}}\;=\;\mathbf{T}_{\text{base}}^{\text{ee}}\;\mathbf{T}_{\text{ee}}^{\text{cam}}\;\mathbf{T}_{\text{cam}}^{\text{pose}},$ (4)
-where:
-  1. 1.
-    $\mathbf{T}_{\text{base}}^{\text{ee}}$ is the forward-kinematics transformation from the robot base to its end-effector,
-  2. 2.
-    $\mathbf{T}_{\text{ee}}^{\text{cam}}$ is the hand–eye calibration matrix (mapping from end-effector to camera),
-  3. 3.
-    $\mathbf{T}_{\text{cam}}^{\text{pose}}$ is the camera-to-pose transformation predicted by our grasp model.
-To execute a grasp in closed loop, we generate a smooth interpolation $\mathcal{C}(t)$ in $\mathrm{SE}(3)$ between the initial end-effector pose $\mathbf{T}_{\text{init}}$ and the desired grasp pose $\mathbf{T}_{\text{goal}}$, as:
-$\mathcal{C}(t)\;:\;[0,\,1]\;\to\;\mathrm{SE}(3),\quad\mathcal{C}(0)=\mathbf{T}_{\text{init}},\quad\mathcal{C}(1)=\mathbf{T}_{\text{goal}}.$ (5)
-At each control step $t$, the desired pose is
-$\mathbf{T}_{\text{des}}\;=\;\mathcal{C}(t).$ (6)
-Let $\mathbf{T}_{\text{current}}$ denote the current pose of the end effector. We define the positional error
-$\mathbf{e}_{p}=\mathbf{p}_{\text{des}}-\mathbf{p}_{\text{current}},$ (7)
-and compute the rotational error using the matrix logarithm of the orientation difference:
-$\mathbf{R}_{\text{err}}=\mathbf{R}_{\text{des}}\,\mathbf{R}_{\text{current}}^{\top},\qquad\boldsymbol{\theta}_{\text{err}}=\mathrm{Log}\!\bigl{(}\mathbf{R}_{\text{err}}\bigr{)},$ (8)
-where $\mathbf{R}_{\text{des}}$ and $\mathbf{R}_{\text{current}}$ are the rotation components of $\mathbf{T}_{\text{des}}$ and $\mathbf{T}_{\text{current}}$, respectively, and $\mathrm{Log}(\cdot)$ is the matrix logarithm on $\mathrm{SO}(3)$.
-We form the combined error vector
-$\mathbf{e}=\begin{bmatrix}\mathbf{e}_{p}\\ \boldsymbol{\theta}_{\text{err}}\end{bmatrix}\;\in\;\mathbb{R}^{6}.$ (9)
-The desired end-effector velocity is then
-$\dot{\mathbf{x}}_{\text{target}}\;=\;K_{p}\;\mathbf{e},$ (10)
-where $K_{p}$ is a diagonal gain matrix. We compute the joint velocity command by projecting $\dot{\mathbf{x}}_{\text{target}}$ through the robot Jacobian pseudoinverse:
-$\dot{\mathbf{q}}_{\text{cmd}}\;=\;\mathbf{J}^{\dagger}\;\dot{\mathbf{x}}_{\text{target}}.$ (11)
-Finally, we update the robot’s joint states with $\dot{\mathbf{q}}_{\text{cmd}}$ and advance $t\!\leftarrow\!t+\Delta t$. This process repeats until
-$\|\mathbf{T}_{\text{des}}-\mathbf{T}_{\text{current}}\|\;\leq\;\epsilon,$ (12)
-indicating successful convergence to the target grasp pose.
-4 RefGraspNet Dataset: Learning to Grasp from 200M+ Grasp Poses
+Most LDRG methods predict an object mask and grasp pose simultaneously.
+A binary object segmentation mask already pinpoints the pixels referenced by the query, implicitly limiting where a valid grasp can lie.
+If the mask is correct, the grasp network need not evaluate locations outside it, shrinking search space and reducing errors.
+We treat the mask as an explicit guiding mechanism: features are pooled within the segmented region before grasp prediction.
+3.3 MapleGrasp
+Given an RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and a referring instruction $T$, we compute vision-language embeddings using a pre-trained CLIP model: $x_{\text{img}}=\text{CLIP}_{\text{img}}(I)$, $x_{\text{text}}=\text{CLIP}_{\text{text}}(T)$.
+The visual embedding $x_{\text{img}}$ is passed through a multi-modal Feature Pyramid Network $\mathcal{F}$, and fused with the text embedding via cross-attention, resulting in $x_{\text{fused}}=\text{CrossAttn}(\mathcal{F}(x_{\text{img}},x_{\text{text}}),x_{\text{text}})$.
+These embeddings are then processed through a series of convolutional and upsampling layers to produce a spatial feature map $f\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$, where $H^{\prime}$ and $W^{\prime}$ are intermediate spatial resolutions and $C$ is the channel dimension.
+We introduce Mask-Guided Feature Pooling within a two-stage training framework to guide grasp prediction.
+In Stage I (Segmentation Training), we freeze the grasp prediction heads and train the segmentation branch independently.
+The text embedding $x_{\text{text}}$ is projected via a linear layer to obtain $z\in\mathbb{R}^{C}$, which is then used to compute the object-centric mask via a per-pixel dot product: $M_{i,j}=\sigma(f_{i,j}^{\top}z)$, yielding a soft binary mask $M\in[0,1]^{H^{\prime}\times W^{\prime}}$.
+This stage allows mild overfitting to ensure accurate object localization for downstream pooling.
+In Stage II (Mask-Guided Grasp Prediction), we train the full model end-to-end.
+The predicted mask $M$ is used to modulate the visual features via element-wise multiplication, producing $f_{\text{pooled}}=f\odot M$.
+These pooled features are refined and passed through three parallel projection heads to predict the grasp quality $Q=\mathcal{H}_{q}(f_{\text{pooled}})$, angle $\theta=\mathcal{H}_{\theta}(f_{\text{pooled}})$, and width $w=\mathcal{H}_{w}(f_{\text{pooled}})$.
+Our goal was to support both 4-DoF and 6-DoF grasping with minimal architectural changes for future dataset compatibility.
+For image-only datasets with 4-DoF grasp pose annotations, we utilize the grasp rectangle to compute ground truth quality, angle, and width maps, following prior work [44].
+For RGB-D datasets with 6-DoF grasp poses, we first identify the 3D area on the target object where ground truth poses occur.
+This region is then projected into 2D to generate a rectangle over the graspable area of the object, which is subsequently treated in a manner analogous to the 4-DoF case.
+MapleGrasp computes the graspable region on the object, which is then overlapped with candidate 6-DoF grasp poses generated by an off-the-shelf grasp sampler, Contact-GraspNet [40].
+The candidate with maximum overlap is selected as the collision-free grasp for the target.
 Dataset Real/Simulated No. of Images Ref. Text Unique Obj. Grasp Poses Obj. Masks Avg. Obj./Img.
@@
 Ours: RefGraspNet 97K 12.25M 88 219M 9.57
-Table 1: Comparison of existing 4-DoF LDRG datasets.
+Table 1: Comparison of existing LDRG datasets.
 RefGraspNet provides 200M+ high quality grasps, object masks for each scene, and thus can be used to finetune both referring segmentation and grasping models.
-We build on existing research and datasets in language driven grasping by creating a new corpus of real world cluttered images and their grasp annotations.
-Our dataset, RefGraspNet, is derived from the GraspNet-1B dataset [11] and comprises a diverse corpus collected using 88 unique objects arranged in cluttered scenes.
-We construct 10 instruction templates such as “Grasp the {object}”, “Find the {object} in this scene”, etc.
-and expand the object names to create referring expressions.
-In scenes with multiple objects, we leverage an open-source model, DeepSeek-VL [30], to provide additional attributes such as color, shape and relative position.
-We chose GraspNet-1B as our base corpus since it consists of 1 billion + grasp poses annotated with sampling and force-closure techniques.
-We sample grasps with a confidence score greater than 70%, and our final corpus comprises 219M grasp poses, 8 times bigger than the previous largest dataset for language driven 4 DoF grasping.
-Comparisons with prior datasets are provided in Tab.
-1.
-We randomly choose 70% of the objects to be seen objects (with the rest as unseen) and create train/val/test-seen and test-unseen splits for our experiments.
+3.4 Loss Function
+Prior work utilizes the Smooth L1 loss for pixel-wise prediction in visual grounding, defined per pixel $(\hat{x}_{p},x_{p})$ as:
+$L_{\mathrm{smoothL1}}(\hat{x}_{p},x_{p})=\begin{cases}\tfrac{(\hat{x}_{p}-x_{p})^{2}}{2\beta},&\!\text{if }|\hat{x}_{p}-x_{p}|<\beta,\\ |\hat{x}_{p}-x_{p}|-\tfrac{\beta}{2},&\!\text{otherwise.}\end{cases}$ (1)
+Here, $\hat{x}_{p}\in\mathbb{R}$ denotes predicted value at pixel $p$, $x_{p}\in\mathbb{R}$ is the ground-truth value, and $\beta>0$ is a constant that determines the transition point between L2 and L1 loss regions.
+We extend this to formulate weighted Smooth L1 loss $w_{p}\,L_{\mathrm{smoothL1}}(\hat{x}_{p},x_{p})$, by assigning per‐pixel weights $w_{p}=1+\alpha\,Q_{\mathrm{gt},p}$, where $Q_{\mathrm{gt},p}\in[0,1]$ is the ground‐truth quality map indicating the object grasp region.
+By emphasizing suitable grasp regions, we enhance precise grasp estimation.
+Our experiments demonstrate that this modified loss function is beneficial in low-resource settings, where the dataset contains only single grasp annotation per object, ensuring more reliable and efficient learning.
+3.5 Metrics
+Language-driven grasp poses are evaluated based on their correctness, feasibility, and overall quality, given a language instruction and the corresponding scene image.
+Quantitative evaluation is performed using grasp success rate for both Top-1 and Top-K predictions.
+For 4-DoF grasping, a prediction is considered successful if its rotation angle differs from the ground-truth by less than $30^{\circ}$, and its intersection-over-union (IoU) with the ground-truth grasp rectangle exceeds 0.25 [31, 45].
+For 6-DoF grasping, success is determined if the selected grasp proposal corresponds to the ground-truth grasp for the referred object [8].
+4 RefGraspNet Dataset: Learning to Grasp from 200M+ Grasp Poses
+To advance language-driven robotic grasping, we introduce RefGraspNet, a richly annotated dataset of over 219 million high-quality 6 DoF grasp poses, derived from real-world cluttered scenes.
+Building on GraspNet-1B [10], which provides over one billion sampled grasps with force-closure scoring comprising 88 distinct household and industrial objects arranged in hundreds of densely cluttered table-top scenarios.
+Each scene captured from multiple calibrated viewpoints yielded object masks and 6 DoF grasp poses.
+We automate generating referring expressions using DeepSeek-VL [28], an open-source VLM.
+Starting from ten instruction templates (e.g., “Grasp the {object},” “Locate the {object}”), we expanded object mentions with color, shape, and spatial qualifiers in multi-object layouts (e.g., “Grasp the red cup behind the yellow bowl”).
+This yielded over 12.25M unique referring phrases, ensuring robust grounding of grasp commands in cluttered environments.
+Grasp candidates were filtered using a 70% force-closure confidence threshold, which was selected based on extensive manual validation, resulting in 219 million validated 6 DoF grasps—roughly eight times larger than the previous largest language-driven grasp dataset.
+We provide comprehensive comparisons in Tab.
+1, demonstrating improvements in scale, scene complexity, and linguistic variety.
+To support empirical studies, we split our objects into 70% “seen” (62 objects) and 30% “unseen” (26 objects), and generate train/val, test-seen, and test-unseen splits.
+This design enables evaluation of generalization to novel objects and arrangements.
+Examples are shown in Fig.
+4.
+Figure 4: Samples from RefGraspNet showcasing cluttered scenes with multiple viewpoints, with diverse objects, distractors, and grasping instructions. Only top-down grasps shown for brevity.
+Model Epochs to Conv. OCID-VLG RefGraspNet
+Test Seen Test Unseen
+DetSeg + CLIP [1] 22 28.12 / 39.21 40.19 / 41.56 27.18 / 27.20
+GR-ConvNet + CLIP [20] 15 9.73 / 15.41 34.19 / 46.56 30.19 / 32.17
+SSG + CLIP [52] 41 33.51 / 34.70 48.78 / 50.10 14.36 / 14.89
+Molmo+SAM2+GraspNet 63.22 / 69.14 68.85 / 70.27 66.23 / 69.41
+CROG [44] 50 77.22 / 87.71 85.32 / 86.49 70.81 / 71.99
+HiFi-CS [2] 44 70.54 / 79.12 73.27 / 74.55 62.13 / 63.28
+ETRG [55] 49 82.28 / 91.12 86.13/88.12 68.11 / 70.05
+Ours: MapleGrasp 32 88.15 / 92.98 89.86 / 91.67 76.92 / 77.67
+Table 2: Benchmarking MapleGrasp across OCID-VLG and RefGraspNet datasets for 4-DoF and 6-DoF vision-language driven grasping respectively.
+Scores reported as grasping accuracy for top-1/top-5 predictions (see Sec.
+3.5).
 5 Experiments
-Here, we present our experiments evaluating MapleGrasp on OCID-VLG and RefGraspNet benchmarks, while also comparing our framework to prior LDRG and VLA methods in simulation and real-world setups.
-5.1 OCID-VLG Benchmark
+We evaluate MapleGrasp on OCID-VLG and RefGraspNet, while comparing it to prior LDRG and VLA methods in both simulation and real-world settings.
+5.1 Benchmark Comparisons
 The OCID-VLG corpus [44] contains 1763 highly cluttered indoor tabletop RGB-D scenes with 31 unique objects.
 Objects have annotated 4-DoF grasp poses and segmentation masks, associated with referring expressions describing attributes such as color, shape, and relative position.
-With 89K unique (RGB-Text-Mask) tuples, it is the largest publicly available benchmark for language-driven grasp detection frameworks [^1] GraspAnything++ is another open-source corpus, but the publicly shared subset is significantly smaller than what was used for training their models, preventing fair benchmarking..
-We evaluate various CLIP-based architectures on OCID-VLG, applying the MapleGrasp framework to Ref2Grab [2], which uses frozen CLIP encoders for improved generalization, and CROG [46], as detailed in Sec.
-3.4.
-Results are summarized in Tab.
-2.
-Model IOU Pr@50 Pr@60 Pr@70 Pr@80 Pr90 J@1 J@Any
-DetSeg + CLIP 29 27.2 20.9 17.5 17.2 16 28.1 39.2
-GR-ConvNet + CLIP 31.3 21.0 11.6 5.5 2.4 0.5 9.73 15.41
-SSG + CLIP 33.6 35.6 35.6 35.5 35.5 32.8 33.5 34.7
-CROG 81.1 96.9 94.8 87.2 64.1 16.4 77.2 87.7
-Ref2Grab 83.26 85.68 85.13 91.53 89.69 83.21 70.5 79.1
-ETRG 80.11 - - - - - 82.28 91.12
-MapleGrasp-Ref2Grab 83.78 85.96 85.41 91.74 90.13 83.79 76.8 84.7
-MapleGrasp-CROG 81.36 97.40 95.32 87.9 65.4 16.4 86.15 91.9
-Table 2: Results on OCID-VLG.
-We implement all baseline methods except ETRG, whose results are taken from their paper.
-Our comparison is limited to methods that do not utilize depth maps during training.
+With 89K unique (RGB-Text-Mask) tuples, it is the largest publicly available dataset for LDRG [^1] GraspAnything++ is open-sourced; however, its public subset is much smaller than the full training data, making fair benchmarking impossible..
+We also provide a benchmark on our new contributed RefGraspNet corpus.
+We evaluate various CLIP-based architectures on both datasets: CROG [44] utilizes CLIP embeddings with cross-attention, followed by MLP layers, to simultaneously predict object masks and grasp maps.
+HiFi-CS [2] employs a frozen CLIP model with FiLM-based fusion and upsampling to generate object masks, followed by grasp estimation using a separate frozen network.
+ETRG [55] introduces trainable adapters within the CLIP model to adapt it for grasping, and is trained on additional data to segment referring expressions and predict referring affordance.
+All baselines are trained for 50 epochs to ensure a fair comparison (see Tab.
+2).
+We also include a strong zero-shot baseline that combines Molmo [6] for identifying the referred object, SAM 2 [37] for generating the object mask, and GraspNet [10] for predicting both 4-DoF and 6-DoF grasp poses.
+MapleGrasp achieves a 7% relative improvement in Top-1 success rate over previous state-of-the-art methods, demonstrating higher precision in grasp prediction in OCID-VLG.
+In RefGraspNet, MapleGrasp outperforms baselines for both seen and unseen objects, with a larger performance margin observed in the unseen setting.
+We attribute this to our closely coupled dual-stage training, which is able to provide partial object masks in unseen objects, which is then used by the grasp detection module to generate graspable object regions and accurate grasp poses.
+MapleGrasp trains more efficiently, converging in 32 epochs ($\sim$8 hours), compared to 45–50 epochs (12–16 hours) required by other baselines on an RTX 4090 GPU.
 Referring Text Molmo+SAM+GraspNet CROG MapleGrasp
@@
 Find the banana furthest away from me [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Pass the tissues to the rear left of the blue and black marker [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Table 3: Qualitative comparisons of LDRG approaches under complex referring queries.
 The zero-shot baseline (first column) struggles to identify referred object due to distractors.
-MapleGrasp (last column) can predict a stable and accessible grasp for each queried object.
-Applying mask proposals enhances both the CROG and Ref2Grab baselines in terms of the Jacquard Index J@N metric, which evaluates the top-N grasp rectangle candidates that exhibit a rotation angle difference of less than $30^{\circ}$ and an IoU greater than 0.25 compared to the ground-truth grasp rectangle.
-MapleGrasp achieves a 12% relative improvement in J@1 over previous state-of-the-art methods, demonstrating higher precision in grasp prediction.
-However, referring expression segmentation (RES) accuracy, measured using Prec@X, which quantifies the percentage of test cases where the IoU exceeds $X\in\{0.5,0.6,0.7,0.8,0.9\}$, remains relatively unchanged.
-This may be due to CROG losing pretrained knowledge during CLIP layer fine-tuning.
-In contrast, MapleGrasp-Ref2Grab, which employs frozen vision-language encoders, achieves higher performance in the RES task but struggles with grasp map prediction due to its limited trainable parameter space (4M vs.
-CROG’s 167M).
-Overall, our findings suggest that MFP enhances both frozen and fully fine-tuned CLIP-based models for language-driven grasp detection.
-5.2 Ablation Studies
-We conduct ablation studies on MapleGrasp (LABEL:tab:ablations) to analyze its effects due to architectural modifications.
-Results highlight the importance of accurate RES for mask pooling; using Molmo, which achieves lower IoU on RES tasks, negatively impacts grasp accuracy due to complex referring queries requiring task-specific visual grounding.
-While using frozen CLIP improves RES, it decreases grasp performance because of fewer trainable parameters.
-Replacing cross-attention with MLP-Mixers, despite reported benefits [44], does not significantly improve performance.
-Furthermore, replacing the weighted smooth L1 loss with a standard Smooth L1 loss leads to reduced scores, likely from slower convergence.
-All ablations use a fixed number of training epochs for consistency.
-Method IoU J@1 J@Any
-MapleGrasp-CROG 81.36 86.15 91.90
-(1) Molmo+SAM for Mask Pooling 73.42 77.79 80.16
-(2) Frozen CLIP Layers 83.52 72.19 74.44
-(3) No Cross-Attn, MLP-Mixer 80.16 84.37 90.45
-(4) Standard Smooth L1 Loss 80.17 83.23 89.18
-5.3 RefGraspNet Benchmark
-We further evaluate previous baselines and MFP-enhanced architectures on our RefGraspNet dataset, establishing a strong benchmark for future research.
-The dataset comprises 88 unique objects, with 70% assigned as seen objects and 30% as unseen objects.
-Results for both test splits are reported in Tab.
-4.
-Similar trends as observed in the OCID-VLG benchmark hold here, with MFP providing significant performance gains for both Ref2Grab and CROG.
-However, all trained models perform worse on unseen objects, highlighting a lack of zero-shot generalization.
-In contrast, open-set object detectors such as Molmo [7] achieve higher test unseen IoU.
-This motivates us to incorporate Molmo in an ensembling setup to boost zero-shot segmentation, which is essential for robust grasp pose estimation.
-Test Seen Test Unseen
-Model IoU J@1 J@Any IoU J@1 J@Any
-DetSeg + CLIP 48.18 40.19 41.56 30.57 27.18 27.20
-GR-ConvNet +
-CLIP 52.34 44.19 46.56 32.18 30.19 32.17
-SSG + CLIP 57.16 48.78 50.10 38.19 34.16 34.89
-CROG 84.49 85.32 86.49 69.95 70.81 71.99
-Ref2Grab 87.32 73.27 75.45 75.67 62.12 63.28
-MapleGrasp-Ref2Grab 87.74 75.67 77.89 75.11 65.19 67.22
-MapleGrasp-CROG 84.86 89.15 89.67 70.43 74.92 75.67
-Table 4: Results on RefGraspNet. The MFP framework leads to consistent improves across seen and unseen objects. However, performance on unseen objects deteriorates for all methods.
+MapleGrasp (last column) can predict an accurate and accessible grasp for queried objects in the first three rows.
+Last row shows an example where all models fail to recognize the correct object (ground truth in green).
+5.2 Qualitative and ablation studies
+We provide comparisons in (Tab.
+3).
+In row 1 and 3, we observe that the off-the-shelf approach fails to accurately predict a grasp on the referred object.
+This occurs because the query contains attributes like “rightmost” and “furthest” increasing the complexity of referring.
+Since the masks are predicted incorrectly, the downstream GraspNet predicts an incorrect grasp.
+While CROG recognizes these objects better, having seen a variety of referring attributes during training, often fails to localize the grasp within the target object.
+MapleGrasp’s mask-guided predictions are accurate in first three rows, but fail for occluded objects and ambiguous queries, where the mask alone cannot localize the grasp pose (row 4).
+More examples in supplementary Sec.
+$\mathcal{S}$2.
+Figure 5: Comparisons across text inputs: from level 1 with just object names to level 4 with color, shape, position attributes
+To compare methods across language complexity, we group test queries by the number of object attributes they mention.
+For example, “Grasp the red circular box near the top right of the image” contains three attributes that an LDRG model must exploit to locate a grasp.
+Using a popular named entity recognition model, GLiNER [57], we count attributes in the RefGraspNet test split and assign queries to four bands—L1 EASY (one attribute) through L4 HARD (four).
+Fig.
+5 plots Top-1 accuracy for MapleGrasp and two baselines: accuracy declines with added attributes, yet MapleGrasp consistently remains ahead.
+We conduct ablation studies on MapleGrasp (Tab.
+4) to analyze its effects due to architectural modifications.
+Results highlight the importance of accurate RES for mask pooling; using Molmo, which achieves lower IoU (50-60%) on object mask detection, negatively impacts grasp accuracy due to complex referring queries requiring task-specific visual grounding.
+While using frozen CLIP improves object identification, it reduces grasp scores due to insufficient grounding of VL features for grasp detection.
+Replacing cross-attention with MLP-Mixers [42], does not improve performance.
+Furthermore, replacing the weighted smooth L1 loss with a standard smooth L1 loss leads to reduced scores due to slower convergence.
+All ablations use a fixed number of training epochs.
+Method Top-1 Top-5
+MapleGrasp 88.15 92.90
+(1) Molmo+SAM for Mask Pooling 78.69 81.27
+(2) Frozen CLIP Layers 73.14 75.57
+(3) No Cross-Attn, MLP-Mixer 85.42 91.54
+(4) Standard Smooth L1 Loss 84.35 90.23
+Table 4: Ablation study on the OCID-VLG corpus.
+Figure 6: Training data efficiency comparisons across baselines on RefGraspNet. Scores reported as grasping success rate in %.
+Lastly, we evaluate the training efficiency of MapleGrasp in comparison to two baseline methods.
+In this experiment, we vary the size of the training dataset and assess model performance after training on 20%, 40%, 60%, 80%, and 100% of the RefGraspNet corpus.
+The results, presented in Fig.
+6, demonstrate that MapleGrasp consistently outperforms the baselines on the test set as the dataset size increases, with particularly notable improvements observed after training on just 20–40% of the data.
+These findings suggest that incorporating referring object masks for pooling vision-language feature maps in grasp detection enhances both training effectiveness and sample efficiency.
 5.3 Cross-Dataset Testing
-To assess the generalizability of models trained on RefGraspNet, we cross-test them on the OCID-VLG benchmark.
-We compare MapleGrasp-Ref2Grab and MapleGrasp-CROG, and present results in Tab.
-5.
-Our findings reveal that models trained on RefGraspNet generalize better than those trained solely on OCID-VLG.
-The dataset’s diversity in grasp poses and text expressions enhances inference on other benchmarks.
-This highlights the advantages of training on RefGraspNet, encouraging further adoption of our corpus to push performance boundaries.
+To assess generalizability of models trained on RefGraspNet, we cross-test them on the OCID-VLG benchmark.
+We compare CROG and MapleGrasp Tab.
+5 and our findings reveal that models trained on RefGraspNet generalize better than those trained solely on OCID-VLG.
 OCID-VLG RefGraspNet
-Model Train-Dataset J@1 J@Any J@1 J@Any
-MapleGrasp-Ref2Gr OCID-VLG 76.8 84.7 41.8 42.9
-RefGraspNet 71.2 74.4 75.7 77.9
-MapleGrasp-CROG OCID-VLG 86.2 91.9 33.2 35.6
-RefGraspNet 68.2 73.8 89.2 89.7
-Table 5: Cross-dataset evaluation results. Models trained on RefGraspNet generalize better than OCID-VLG.
-5.5 LIBERO Simulation Experiments
-We deploy the MapleGrasp model trained on RefGraspNet and extend grasp poses to 3D using an inverse kinematics controller for trajectory prediction (Sec.
-3.6).
-This setup allows direct comparisons with Vision-Language-Action (VLA) models.
-Since our approach focuses on language-driven grasp pose estimation, we cannot evaluate task success rates directly.
-Instead, we define grasp success as trials in which an object is lifted and held above ground level for at least two seconds.
-We report J@1 and J@Any metrics.
+Model Train-Dataset Top-1 Top-5 Top-1 Top-5
+CROG OCID-VLG 77.2 87.7 41.8 42.9
+RefGraspNet 68.2 73.8 85.32 86.49
+MapleGrasp OCID-VLG 88.2 92.9 43.2 45.6
+RefGraspNet 79.5 81.7 89.2 89.7
+Table 5: Cross-dataset evaluation results (highlighted). Models trained on RefGraspNet generalize better than OCID-VLG.
+5.4 LIBERO Physics Simulation Experiments
+We test MapleGrasp in the LIBERO simulation environment in a table-top setting.
+The grasp poses generated by MapleGrasp are executed with an inverse kinematics velocity controller described in supplementary Sec.
+$\mathcal{S}$3.
+We compare our modular grasping+trajectory optimization method with end-to-end VLAs.
+While VLAs are trained on tele-operated demonstrations of successful tasks, we use these observations to identify stable grasp poses for referred objects.
+We only benchmark the object grasping skill.
+For example, consider the task “Put the wine bottle on the rack.” We consider the task to be successful if the robot grasps the “wine bottle” and lift it above the ground.
+Success rate is measured manually and averaged with 10 trial runs per task.
+Model SPATIAL GOAL OBJECT
+Diffusion Policy 77 69 92
+Octo 78 84 88
+OpenVLA 84 79 90
+MapleGrasp 87 85 90
+Table 6: Libero simulation results across SPATIAL, GOAL and OBJECT task suites. Scores are reported using success rate
 Experiments are conducted on three task suites from the LIBERO benchmark: SPATIAL, GOAL, and OBJECT.
-Since VLAs are fine-tuned on 50 teleoperated demonstrations per task, we automate 4-DoF grasp annotations using forward kinematics-based 3D grasp pose estimation and project the masks to 2D, inverting steps described in Sec.
-3.1.
-We fine-tuned MapleGrasp on these annotations to adapt to the simulation environment.
-At inference time, MapleGrasp predicts 2D grasps, which are transformed and executed in 3D by the controller.
+For more details about the tasks, refer to supplementary Sec.
+$\mathcal{S}$4.
 Results in Tab.
 6 compare our method with VLA and diffusion baselines.
-Model SPATIAL GOAL OBJECT
-Diffusion Policy 0.77 0.69 0.92
-Octo 0.78 0.84 0.88
-OpenVLA 0.84 0.79 0.90
-MapleGrasp-CROG 0.87 0.85 0.90
-Table 6: Libero simulation results across SPATIAL, GOAL and OBJECT task suites. Scores are reported using J@1.
 Our results demonstrate that our grasping+control pipeline achieves similar accuracy to recent VLAs in grasping tasks.
@@
 Model Train-Dataset GOAL OBJECT
-OpenVLA GOAL 0.79 0.00
-OBJECT 0.00 0.90
-MapleGrasp-CROG GOAL 0.85 0.62
-OBJECT 0.68 0.90
-Table 7: Comparison of OpenVLA and our method in cross-dataset settings. OpenVLA struggles with seen objects under novel tasks, while our method generalizes well to unseen queries.
+OpenVLA GOAL 79 0
+OBJECT 12 90
+MapleGrasp GOAL 85 62
+OBJECT 68 90
+Table 7: Cross-dataset results show OpenVLA struggles on seen objects in new tasks, while MapleGrasp generalizes better.
 5.5 Real-Robot Experiments
-Figure 3: Setup for robot trials using 7 DoF franka arm and our modular pipeline described in section Sec. 3.6
-We evaluate various LDRG methods in a real-world robotic setup using a Franka arm with a mounted camera at the gripper, as shown in Fig.
-3.
-We construct diverse tabletop scenes featuring single and multiple objects, distractors, and varying lighting conditions.
+Figure 7: Setup for robot trials using 7 DoF franka arm.
+We evaluate various LDRG methods in a real-world robotic setup using a Franka arm with a camera mounted at the gripper, as shown in Fig.
+7.
+All models were fine-tuned on a combined dataset of RefGraspNet and OCID-VLG.
+For evaluation, we collected 15 objects, 10 seen and 5 unseen to create randomly arranged tabletop scenes (see supplementary Sec.
+$\mathcal{S}$5).
+To test capabilities of all baselines, we performed separate experiments trying to grasp seen and unseen objects in the presence of distractors (objects that look similar to the target object either in shape, color, or type).
+Each grasping instruction was tested 10 times and averaged.
+This resulted in N = 300 (15x2x10) robot trials.
 Results are presented in Tab.
 8.
-For real-robot trials, MapleGrasp integrates Molmo-based object mask refinement, which compares our model’s segmentation predictions with Molmo’s out-of-vocabulary predictions, selecting the highest mutual overlap.
-MapleGrasp outperforms other baselines, achieving a notable 7% relative improvement in cluttered setups.
-Model Single Cluttered
-Det-Seg+CLIP 0.27 0.21
-GR-ConvNet+CLIP 0.30 0.27
-Molmo+SAM+GraspNet 0.45 0.47
-SSG+CLIP 0.21 0.17
-Ref2Grab 0.42 0.40
-CROG 0.55 0.53
-MapleGrasp-CROG 0.59 0.57
+We consider a trial to be successful if the robot grasps object and lifts it 30cm above the surface.
+[…]
+Seen Objects Unseen objects
+Model Single Cluttered Single Cluttered
+Molmo+SAM2+GraspNet 70 61 71 66
+HiFi-CS 72 63 65 59
+CROG 85 73 69 54
+MapleGrasp 93 87 72 65
+MapleGrasp+Molmo 91 85 74 73
 Table 8: Real robot experiments. Success rate measured across 10 trials for every grasp instruction.